40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 27 01:58:23 2011, kiwamu, Update, LSC, locking status 
    Reply  Wed Jul 27 08:59:10 2011, Suresh, Update, LSC, locking status MC_RFPD_DCMon.png
    Reply  Wed Jul 27 10:10:12 2011, steve, Update, PSL, locking status 80dpmcMC.jpg
Message ID: 5040     Entry time: Wed Jul 27 01:58:23 2011     Reply to this: 5041   5043
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: locking status 

Through some locking exercise I found that several things are degrading.

Remember the interferometer is like a cat, so we have to feed and take care of her everyday. (Otherwise the cat will be dead !)

 

Beam axis:

 I guess that the beam axis has changed a lot to the horizontal direction.
The beam spots on the REFL and AS camera looked off-centered by a size of the spot.
The beam axis has to be well-aligned before the vent.

 

Locking of the Arms :

 didn't lock at all. It could be a problem of the demodulation phase on AS55.
Also the TRY camera looked pretty much off-centered. The spot is already getting out from the field of view.
We have to fix this issue, otherwise we cannot align the beam axis.
 
 

Locking of PRM :

 Sort of okay, I was able to lock both MICH and PRCL although I had to flip the sign of the MICH control gain due to the demod-phase change.
The suspensions don't look healthy. The beam spots on the REFL and AS camera move a lot even without any length feedback.
It means some of the suspensions are shaky.
ELOG V3.1.3-